# Learning Generic and Dynamic Locomotion of Humanoids Across Discrete Terrains

### v1

@@ II-A Contact Implicit Trajectory Optimization (CI-TO)
Contact, because of its binary nature, introduces discrete jumps in the gradient when it is incorporated into the optimization variables. Commonly used techniques, such as mixed-integer programming [43] and linear complementary problem (LCP) [44], share this foundational difficulty. Consequently, real-time control using CI-TO optimization needs to tackle two major issues: prolonged computation times and the risk of converging to local minima that yield spurious results. [45] suggested a potential solution by introducing soft contact constraints. More recently, advances in differential dynamic programming (DDP)-based approaches have demonstrated experimental success with quadruped robots [20, 18]; however, these require the optimization to start near the solution due to DDP’s local search nature. Alternatively, [19] addressed the real-time landing contact sequence problem by solving LCP with a pre-trained warm starting point generator. This approach effectively reduced computation time but showing limited performance beyond the trained data set. [46] circumvented the gradient discontinuity issue by representing the contact sequence with parameterized polynomial functions. However, this method necessitates specifying the number of contact points and has not been tested in full dynamics simulations.
@@ III Basal Ganglia Policy Training
where $\mathbf{p}\in\mathbb{R}^{2}$ and $\mathbf{v}\in\mathbb{R}^{2}$ are the rigid body’s center of mass position and velocity, respectively, with respect to the stance foot frame. $\theta\in\mathbb{R}$ and $\omega\in\mathbb{R}$ are the pitch angle and angular velocity, respectively, and $\mathbf{h_{t}}\in\mathbb{R}^{20}$ is the height map of the terrain (Fig. 2(a)). Our BG-policy makes an action when the right foot is under contact, where the origin of the local frame is attached to. The policy outputs the desired forward velocity, the contact locations of the next two steps (left foot and right foot) in the local frame, and the gait type (Fig. 2(b)). Subsequently, a trajectory optimisation seeks the optimal trajectory for the upcoming two steps (Fig. 2(c)) considering the contact constraints set by given action along with the current state and other constraints (e.g., reaction force and kinematics limit). The last state of the optimal trajectory, with some random Gaussian noise added, returns to the policy as the next state of the environment (Fig. 2(d)).
The state of the planar single rigid body model is given by

### v2

@@ II-A Contact Implicit Trajectory Optimization (CI-TO)
Contact, because of its binary nature, introduces discrete jumps in the gradient when incorporated into optimization variables. Common techniques like mixed-integer programming (MIP) [43] is suffered from exponentially growing computational time. On the other hand, contact implicit trajectory optimization (CI-TO), as initially presented in [44] and later in [45], use complementarity constraints to mitigate this issue. However, real-time control using CI-TO faces the risk of converging to local minima. Soft contact constraints have been suggested as a solution  [46]. More recently, advances in differential dynamic programming (DDP)-based approaches have demonstrated experimental success with quadruped robots but require starting near the solution due to DDP’s local search nature  [20, 18]. Alternatively, solving LCP with a pre-trained warm starting point generator has reduced computation time but performed poorly beyond the trained data set  [19]. [47] circumvented the gradient discontinuity issue by representing the contact sequence with parameterized polynomial functions. However, this method necessitates specifying the number of contact points and has not been tested in full dynamics simulations.
@@ III Basal Ganglia Policy Training
where $\mathbf{p}\in\mathbb{R}^{2}$ and $\mathbf{v}\in\mathbb{R}^{2}$ are the rigid body’s center of mass position and velocity, respectively, with respect to the stance foot frame. $\theta\in\mathbb{R}$ and $\omega\in\mathbb{R}$ are the pitch angle and angular velocity, respectively, and $\mathbf{h_{t}}\in\mathbb{R}^{20}$ is the height map of the terrain (Fig. 2(a)). Our BG-policy makes an action when the right foot is under contact, where the origin of the local frame is attached to. The policy outputs the desired forward velocity, the contact locations of the next two steps (left foot and right foot) in the local frame, and the gait type (Fig. 2(b)). Subsequently, a trajectory optimisation seeks the optimal trajectory for the upcoming two steps (Fig. 2(c)) considering the contact constraints set by given action along with the current state and other constraints (e.g., reaction force and kinematics limit). The last state of the optimal trajectory, with some random Gaussian noise added, returns to the policy as the next state of the environment (Fig. 2(d)). The state of the planar single rigid body model is given by